OpenStudy (anonymous):

simplify the expression. (15m^3 n^-2 p^-1/25m^-2 n^-4)^3

OpenStudy (anonymous):

\[\frac{3375 m^3}{n^{18} p^{3/25}} \]

OpenStudy (anonymous):

can you show how you got to the answer

OpenStudy (anonymous):

From Mathematica 8 Home Edition:\[(15m{}^{\wedge}3 n{}^{\wedge}-2 p{}^{\wedge}-1/25m{}^{\wedge}-2 n{}^{\wedge}-4){}^{\wedge}3=\frac{27 m^3}{125 n^{18} p^3} \]With parenteses added to designate the implied order of grouping and evaluation:\[\text{ }(\text{ }(15m{}^{\wedge}3 )(n{}^{\wedge}-2)( p{}^{\wedge}-1)/25(m{}^{\wedge}-2)( n{}^{\wedge}-4)\text{ }){}^{\wedge}3=\frac{27 m^3}{125 n^{18} p^3} \]From WolframAlpha.com:\[\text{ ( (15m${}^{\wedge}$3 )(n${}^{\wedge}$-2)( p${}^{\wedge}$-1)/25(m${}^{\wedge}$-2)( n${}^{\wedge}$-4) )${}^{\wedge}$3 =} \frac{27 m^3}{125 n^{18} p^3} \]WolframAlpha.com, no parentheses:\[\text{(15 m${}^{\wedge}$3 n${}^{\wedge}$-2 p${}^{\wedge}$-1/25 m${}^{\wedge}$-2 n${}^{\wedge}$-4)${}^{\wedge}$3 = } \frac{3375 m^3}{n^{18} p^{3/25}} \]There appears to be a discrepancy between the two sites. In general worlfram takes the view that n^a*b is (n^a)*b . If ab is the exponent, they want to see n^(ab) explicitly. Will post a response on this tread regarding this matter if Wolfram responds.

OpenStudy (anonymous):

sorry. thread, not tread